# ProACT: An Augmented Reality Testbed for Intelligent Prosthetic Arms

### v1

@@ 2.4 Study design
Eight right-handed non-amputees (ages $22$-303030305555M, $3$F) gave informed consent for a within-participants study under the oversight of an Institutional Review Board [^3] Protocol 65022, approved by the Administrative Panels for the Protection of Human Subjects at Stanford University, on February 29, 2024.. Each participant attempted $3$ trials of the modified BBT described in Section 2.2 with each of the $4$ methods. The sequence was counterbalanced (Table 3), alternating direct (A, B) and assisted (C, D) methods, so that the assisted methods, with similar participant controls, were never used contiguously.
Table 3: Counterbalanced sequence of methods
@@ 8 C A  D*  B*
Protocol: After filling a short survey regarding their prior experience with human-robot interaction devices, participants received a standardized verbal description of the task. They trained and tested the EMG classifier using a simple custom game shown on a screen. Next, participants wore the HMD, and were seated on a chair, centering it relative to the box and as close to it as possible without intersection. They then received demonstrations for methods A, B and C. For A and B, they were allowed to try moving and switching through each DoF, as well as toggling the hand state, until they understood the operation conceptually [^4] Based on pilot studies which showed that participants found them tedious and confusing at first, performing a full pick and place with these methods would have extended the demonstration phase of the study to unpredictable lengths.. For C, since instructions were dependent on the stages of the task, they were guided verbally through one pick and place. EMG training was repeated if necessary, before beginning trials. A $5$-minute countdown timer was shown on a laptop screen placed in front. Participants were asked to prioritize completion rate (maximize number of blocks transferred before timeout), speed (minimize total time taken), and accuracy (minimize distance between actual and desired placement location), in that order. EMG training was repeated as necessary, if it deteriorated during trials.
Surveys: The commonly-used NASA-TLX [61] was avoided, considering that its length and repetition would encourage satisficing [62] in a tiring task [63, 64]. Instead, a custom $5$-point, $3$ question survey was given, querying the perceived ease, sense of agency and speed of performing the task. At the end of the experiment, participants ranked the methods based on the same questions and were asked which method they would choose if they could “play one last game with any one method”, and the reason for their choice.
@@ 3 Results
Analyses were planned for: (1) block outcomes (success rate in reaching targets), (2) pick duration and place duration when successful, and (3) placement accuracy (minimum distance achieved from target position over the course of the trial). Block outcomes dominated the results, with C and D outperforming A and B uniformly across participants. Due to the high failure rates, only sparse data were available from A and B for the remaining metrics. Therefore, quantitative results are presented for success rates, and the remaining results reported through visualizations.
Figure 4: The outcome shown for each individual block in the experiment.
Figure 5: Pooling participant data, the number of blocks for each outcome, shown sorted by method. $96$ blocks were presented with each method. The percent of blocks that successfully reached their targets is noted for each method.
3.1 Success
Each participant performed $12$ trials, and therefore had the opportunity to interact with $48$ blocks in total. Each block among these had exactly one of the outcomes shown in Fig. (4). There was considerable variation in performance across participants. We can make the following observations: First, from Fig. (4), the two methods with the highest number of successful transfers (out of $3\times 4$) are C and D, for every individual participant. Second, most participants had difficulty using B at first ($5$ of them had no success at all in the first trial with B), but improved in subsequent trials. Third, in the aggregated data from all participants (Table 4), the mean and median number of successes per trial (over $8\times 3$ trials, with $4$ being the maximum success rate per trial) are both higher with C and D. The range and standard deviation are both smaller. Thus, C and D were more successful on average, and more uniformly so, even across participants. Finally, Fig. (5) shows the distribution of outcomes within each method, pooling all the blocks moved by all participants. For every failure outcome, fewer blocks had that outcome with C, D than with A, B. Method B had the highest number of blocks which crossed the partition without reaching the target. The most blocks were dropped on the floor with A.
Table 4: Trial success rates aggregated over participants.
@@ 3.1 Success
3.2 Transfer timings and accuracy
In every trial, participants had the freedom to operate the arm while moving its base (their shoulder), or to simply move the base without actively using the method. The pick and place actions were annotated with this information, based on the joint angle trajectories of the MPL.
Pick duration was defined as the time from the release of the previous block (or start of the trial) until contact was made. Figure. (7(a)), shows the time taken for those picks which were stable enough for the block to eventually cross the partition ($266$ in number; although, for blocks grasped multiple times, the first time was used). The duration of a successful place action was defined as the time from grasping to release, under the conditions that (a) the block crossed the partition during this period, and (b) the target region (within $0.05$ m of the target) was entered before the block was released. These durations ($258$ in number) are shown in Fig. (7(b)). Finally, for those blocks which successfully reached the target regions ($266$ in number, including those dropped slightly outside and pushed), the minimum distance achieved from the target is plotted in Fig. (8).
All three plots reflect the stark disparity in success rates. Figure (7(a)) shows that with A and B, it was faster not to operate the arm than to attempt picking using the method, while C and D enabled picking of many more blocks at comparable rates. For placing, B, C and D share a common trend of participants opting not to use the method slightly more often than to use it. The accuracy of placing [^5] Blocks which crossed the partition but did not reach the target (more often in A and B), are excluded here., conditioned on success, shows no additional difference between methods (Fig. (8)). This is unsurprising, as accuracy was given lowest priority, and the target regions were displayed (Fig. (1b)).
3.3 Questionnaire responses
Participants do not seem to have perceived any significant difference between C and D, but unambiguously preferred them over A and B (Fig (6), Table 5).
Figure 6: Survey rating results after each method.
Table 5: Final survey responses ranking methods.
@@ 8 D $>$ C $>$ A $>$ B D $>$ C $>$ A $>$ B D $>$ C $>$ A $>$ B D
Figure 7: (a) Pick and (b) place durations, excluding extreme outliers (lower is better; details in Section 3.2). The numbers of actions performed without/with each provided method are noted along the “Method” axis.
Figure 8: Average of the minimum distance of block from target. The numbers of place actions performed without/with each provided method are noted along the “Method” axis.
@@ 4 Discussion
The large variation among participants seen in the above results is a typical feature of user studies with assistive technology, which is encouraging to see in this new platform. The excellent performance of Participant 1 (P1) across all methods demonstrates that the task was designed to be possible to complete within the allotted time; at the same time, the remaining participants exhibit a wide range of skill: P8 had a relatively poor success rate, close to $50\%$, across methods, while P2 and P5 showed a large improvement in performance between methods A and D (Fig. (4)).
@@
The testbed highlights the strengths of the new methods – pick times are reduced and made more consistent – which allowed participants to achieve higher success rates, while place times were not improved, and in some cases, actually increased. Anecdotally, this can be explained by the specific implementation of the placing method in C and D. Multiple participants reported difficulty in making the gaze-tracking marker appear on the target side, during the first placing action of trials. If they either succeeded, or decided to use body movement instead, then subsequent transfers were completed quickly. By using the testbed to highlight this failure of the methods being prototyped, we are able to identify an area of future work. We are also able to identify that the primary bottleneck for successful transfers in this task is not placing, but reaching to pick, which requires more precise control (Fig. (7)), explaining why the methods which improve picking lead to unambiguously higher success rates.
@@
It is for a similar reason that the differences between C and D are not obvious from this study; reaching to pick, and not selecting the object to be picked, was the major challenge, which both methods alleviated similarly. In future work, we shall investigate cluttered scenes, in which mere proximity of the center of gaze to different objects might be insufficient.
@@
The nature of the BBT allowed P1 to achieve very fast transfers through skilled movement of the upper body as a whole. While all participants were informed that they were free to move thus, and in fact, would need to do so in order to reach all blocks, only one was able to use body movement to such an extent and with such skill. Had all participants adopted this strategy, no significant results might have been obtained. This motivates future studies based on a more spatially complex task, where objects would not lie on the same plane, and the intended direction of their displacement would not be along a single axis.
The small sample size of this study and the disparate success rates precluded statistical analyses of performance metrics and survey responses, but results suggest that, given the notable random effect of participant skill, mixed-effects models would be most appropriate for larger datasets in the future, with Friedman tests for the pseudo-quantitative components.
@@
No deception was used in this study; participants were possibly biased in favor of C and D, perhaps slightly favoring D. It is difficult to judge whether the minor improvement with D relative to C in success rates may be attributed to any assumed superiority. However, as D occurs before and after C the same number of times, learning and fatigue can both be ruled out as explanations. The effect of modeling sequential tasks merits further investigation with a richer dataset.
@@
Six of eight participants selected D as the method they would most prefer if attempting an additional trial. P2 selected C, which is explained by the deterioration of EMG acquisition in the first iteration of D (Table 3). P7 selected B, explaining that, even though it was more difficult than C/D, it made him feel the most satisfied on success, because he felt in control. This is especially interesting in light of the fact that during trial with B, P7 had declared a general weakness in spatial reasoning, which was most challenged by end-effector control. Understandably, P7 alone ranked B above D on the question of agency. P1 ranked A as second in agency and perceived speed, which may be explained by the technical issue in the first iteration of C. Aside from these, all participants ranked C and D as the top two in all three metrics: ease, agency, and perceived speed. The choice to move the arm along the planned trajectory only as long as EMG input was actively provided seems to have been sufficient to provide a sense of agency to most participants.
@@
More interestingly, in the comparison between A and B, exactly half of the participants preferred A over B and vice-versa. At first glance, this may appear to show some natural variation in preferences. However, examining Tables (35) in conjunction reveals that each participant preferred whichever of these two methods they had performed first; very likely a fatigue effect. It is encouraging that in the light of such a strong effect, the improvements offered by C and D (which were encountered in alternation, as shown in Table 3) overwhelmed the effect of fatigue to keep them at the top of nearly all rankings.
@@ 5 Conclusion
While considerable future work with greater powered samples, as well as amputee populations, is required in the future, the present work demonstrates the following:
@@
    Gaze-based intent estimation and low-level autonomy improve the performance and comfort of users of high-DoF prosthetic arms in complex manipulation tasks.
@@
Due to the open-source nature of much of the latest robotics infrastructure, as well as ProACT itself, it is easy to incorporate more sophisticated methods of intent prediction, and test them in the context of more challenging tasks, a different level of limb loss, or other wearable robotic manipulators.

### v2

@@ 2.4 Study design
The modified BBT described in Section 2.2 was used as a basis for two user studies, to test the reliability of the platform and gain preliminary insights into the methods. Participants gave informed consent under the oversight of an Institutional Review Board [^5] Protocol 65022, approved by the Administrative Panels for the Protection of Human Subjects at Stanford University, on February 29, 2024 and were all right-handed non-amputees.
In Study 1, $8$ participants (ages $22$-303030305555M/$3$F) used all $4$ methods for $3$ trials each, within a single $120$-minute session. The method sequence was counterbalanced using a partial Latin square (Table 3), to evenly distribute the effects of skill transfer and fatigue. Direct (A, B) and assisted (C, D) methods were presented alternately, so that the assisted methods, with similar participant controls, were never used contiguously.
In Study 2, methods A and D were examined at greater length, with $6$ male participants (ages $22$-31313131666P1 from Study 1 volunteered as P2 in Study 2 and was given all the same instructions as other participants. Other participants were naïve.) performing $12$ trials with a single method in each of two $120$-minute sessions ($1$-$4$ days apart). The goal was to collect richer data per method, as high failure rates had impacted quantitative data collection in the shorter Study 1, and to alleviate fatigue. In order to get sufficient data within a limited duration, all participants were administered the methods in the sequence D-A, while recognizing that a skill transfer effect might be seen on the second day. This decision was based on the relatively greater difficulty seen with method A in Study 1, when participants were not given enough time to acclimate to the task.
Together, the two studies show how ProACT can be used to (1) quickly compare and prototype multiple interfaces for intelligent prostheses, and (2) perform detailed investigations on parameters relating to movement and task completion.
Table 3: Counterbalanced sequence of methods in Study 1
@@ 8 C A  D*  B*
Protocol: After filling a short survey regarding their prior experience with human-robot interaction devices, participants received a standardized verbal description of the task. They followed a screen-guided training procedure for EMG gesture classification, holding each gesture for $3$ seconds and performing the sequence of gestures twice, while seated with horizontal unsupported forearm (as it would be held during the trials). Classification performance was evaluated using an on-screen cursor control test. During this familiarization period, participants also had a chance to adapt their gestures for clarity. In Study 1, the protocol was continued when participant and experimenter were both satisfied with classifier performance. In Study 2, participants were asked (at a minimum) to perform all gestures in a prescribed sequence, three times consecutively, without misclassification. Next, participants wore the HMD, performed gaze calibration, and were seated on a chair for demonstrations. In Study 1, for A and B, they were allowed to try moving and switching through all DoF, as well as toggling the hand state, until they understood the operation conceptually [^7] Based on pilot studies which showed that participants found them tedious and confusing at first, performing a full pick and place with these methods would have unpredictably extended the demonstration phase of the study.. For C, since instructions were dependent on the stages of the task, they were guided verbally through one pick and place. If necessary, EMG training was repeated , and a $5$-minute break given, before beginning trials. In Study 2, due to the additional time available, they were allowed to complete two block transfers with each method (D/A) as practice, followed by a $5$-minute break, and with another $2$-minute break after $6$ of $12$ trials [^8] In Study 1, natural breaks from the task occurred during the intermediate surveys. Additional breaks were given if requested, but were not enforced.. In Study 1, a $5$-minute countdown timer was shown on a laptop screen placed in front, which was omitted in Study 2 so as to obtain gaze tracking data pertinent to the task alone. Participants in both studies were asked to prioritize completion rate (maximize number of blocks transferred before timeout), speed (minimize total time taken), and accuracy (minimize distance between actual and desired placement location), in that order. Importantly, in Study 1, they were explicitly told that they were free to move their upper body, while in Study 2, they were asked to limit such motion, actively trying to use the provided methods except when necessitated by reachability constraints. Trials were monitored through Gazebo and RViz rather than streaming the participant’s view, so as not to impede communication and rendering on the HMD.
Surveys: In Study 1, long questionnaires were avoided, to discourage “satisficing” [65] due to time pressure and fatigue [66, 67]. Instead, a custom $5$-point, $3$ question survey was given, querying the perceived levels of ease, sense of agency and speed. At the end of the experiment, participants ranked the methods based on the same questions and were asked which method they would choose if they could “do one last trial with any one method”, with the reason for their choice. In Study 2, which was split into separate sessions and allowed for more time, the Pros-TLX instrument [68] (kindly brought to our attention by a reviewer) was used instead, and subjective comments noted.
@@ 3 Results
Selected results of interest from the two studies are presented below, with the complete data available on the project website.
3.1 Study 1
Block outcomes (success rate in reaching targets) dominated the results, with C and D outperforming A and B uniformly across participants.
Figure 4: The outcome shown for each individual block in the experiment (Study 1).
3.1.1 Success
Each participant performed $12$ trials, and thus had the opportunity to interact with $48$ blocks in total. Each block among these had exactly one of the outcomes shown in Fig. (4). There was considerable variation in performance across participants. We can make the following observations: First, the two methods with the highest number of successful transfers (out of $3\times 4$) are C and D, for every individual participant. Second, most participants had difficulty using B at first ($5$ of them had no success at all in the first trial with B), but improved in subsequent trials. Third, in the aggregated data from all participants (Table 4), the mean and median number of successes per trial (over $8\times 3$ trials, with $4$ being the maximum success rate per trial) are both higher with C and D. The range and standard deviation are both smaller. Thus, C and D were more successful on average, and more uniformly so, even across participants. Finally, for every individual failure outcome, fewer blocks had that outcome with C, D than with A, B. Method B had the highest number of blocks which crossed the partition without reaching the target. The most blocks were dropped on the floor with A.
Table 4: Trial success rates aggregated over participants (Study 1).
@@ 3.1.1 Success
3.1.2 Transfer timings and accuracy
In every trial, participants had the freedom to operate the arm while moving its base (their shoulder), or to simply move the base without actively using the method. The pick and place actions were annotated with this information, based on the joint angle trajectories of the MPL. Pick duration was defined as the time from the release of the previous block (or start of the trial) until contact was made. Figure. (6(a)), shows the time taken to perform those picking actions which were stable enough for the block to eventually cross the partition ($266$ in number; although, for blocks grasped multiple times, the first time was used). The duration of a successful place action was defined as the time from grasping to release, under the conditions that (a) the block crossed the partition during this period, and (b) the target region (within $0.05$ m, or one block diameter, of the target) was entered before the block was released. These durations ($258$ in number) are shown in Fig. (6(b)).
Both plots reflect the disparity in success rates. Figure (6(a)) suggests that with A and B, it was faster not to operate the arm than to attempt picking using the method, while C and D enabled picking of many more blocks at comparable rates. For placing, B, C and D share a common trend of participants opting not to use the method slightly more often than to use it. The accuracy of placing, conditioned on success, shows no additional difference between methods.
3.1.3 Questionnaire responses
Participants do not seem to have perceived any significant difference between C and D, but preferred them over A and B (Fig (5), Table 5).
Figure 5: Survey rating results after each method (Study 1).
Table 5: Final survey responses ranking methods (Study 1). Identical answers are shaded in the same color.
@@ 8 D $>$ C $>$ A $>$ B D $>$ C $>$ A $>$ B D $>$ C $>$ A $>$ B D
Figure 6: Pick and place durations from Study 1, excluding extreme outliers (lower is better; details in Section 3.1.2). The numbers of actions performed without/with each provided method are noted along the “Method” axis.
3.2 Study 2
Study 1 suggests that the two assisted methods might be conducive to performance when participants are first introduced to each method, but it does not capture the learning effect expected as participants gain more experience with each method. It also shows a clear effect of fatigue, albeit averaged across participants through counterbalancing. Therefore, Study 2 further examined performance with the two “extremes” (A and D) in greater detail.
Participants 1, 2, and 6 reported having some prior experience with EMG controllers. It was noted that P3 and P5 complained about EMG misclassification throughout method D, with P3 taking about $13$ minutes to retrain the classifier after trial 2. Data from these sessions were retained, as they show the effect of classification uncertainty on behavior and performance. Participant 4 also retrained the classifier after trial 2, but did not complain during subsequent trials [^9] Note: P4 later reported that until trial 8, he had not internalized that the wrist needed to be held extended for executing motion plans..
During their respective second sessions, all three (P3, P4, P5) were able to achieve highly satisfactory classification (better than any previous instance) within $1$-$3$ attempts. Results show a split between participants whose performance seemed to benefit from method D, and those who performed nearly the same with both methods. For clarity, those who show similar trends (P2, P4, P6) are referred to as Group 1 below. The members of Group 2 (P1, P3, P5) are addressed individually.
3.2.1 Success rates
The total numbers of successful transfers per trial, progressing across trials, are shown in Fig. 7. Group 1 participants showed higher success rates with D than with A.
Figure 7: Success rates across trials (Study 2); participants 2, 4, and 6 did consistently better with method D, while others’ performance fluctuated.
3.2.2 Transfer timings and accuracy
As participants were explicitly asked to use the provided methods as much as possible, we did not make similar annotations as in Fig. 6(a). Group 1 seem to have performed pick actions somewhat faster and more consistently with Method D (Fig. 8(a)), as well as placing blocks with somewhat higher accuracy and greater consistency, including that fewer blocks were disturbed after having entered the target region (Fig. 8(b)). Group 2 seem to perform about the same with either method.
Figure 8: Detailed plots showing speed and compensatory motion during pick and place stages (Study 2).
3.2.3 Compensatory movement
The cumulative motion of the shoulder coordinate frame was calculated for each successful pick and place. Rotation was quantified by computing the angle of rotation (about the appropriate axis) in each time step at $10$ Hz and summing up these angles; therefore, its absolute value, which depends upon the time step, is not meaningful, but the relative values can be compared. In Fig. 8(d), Group 1 appear to have used the rotation of the shoulder to a greater extent when using method A. Perhaps because translation is easier to notice and avoid, the translation data do not show a similar trend (Fig. 8(c)). The slightly greater rotation by P5 and translation by P1 in A may partially explain their comparable performance between A and D along the above metrics.
3.2.4 Gaze
Between the two methods, participants spent a similar fraction of the total transfer time looking at the arm (including mode display). Within the parts of the arm, it is interesting to note that for all participants, more or less consistently across trials, visual attention was more often directed towards the hand in D when compared to A. Additionally, it was observed that (likely because D includes stages when gaze is actively engaged) the fraction of time spent looking at place targets is greater with D.
3.2.5 Questionnaire responses
Weighted scores from Pros-TLX show a mix of results. Conscious processing and visual attention required were perceived to be higher in A by most participants, with uncertainty being higher in D. Overall, all participants except P3 perceived higher workload in A.
@@ 4 Discussion
Results from both studies show that, at least for a section of the (small) sample studied, gaze-based assistance can provide a notable improvement in performance. More importantly, the proposed platform not only also illuminates the large variation among participants that is typical of user studies with assistive technology, but allows us to examine factors which led to improvement or degradation in performance – whether for an individual participant, individual trial, or individual object interaction – and to iteratively design better control strategies.
@@
The ability to track compensatory motions and isolate actions relying upon them is a valuable tool in design. For example, Fig. 6(b) made it clear that C and D were not helpful to participants in the place stage. Anecdotally, this could be explained by the specific implementation of the placing method. Multiple participants reported difficulty in making the gaze-tracking marker appear on the target side, during the first placing action of trials. If they either succeeded, or decided to use body movement instead, then subsequent transfers were completed quickly. By using the testbed to highlight this flaw, we were able to correct it before Study 2, by preemptively displaying the marker. We were also able to identify that the primary bottleneck for successful transfers in this task was not placing, but reaching to pick, which requires more precise control (Fig. (6)), explaining why the methods which improve picking might have led to higher success rates.
@@
It is for a similar reason that the differences between C and D are not obvious from this study; reaching to pick, and not selecting the object to be picked, was the major challenge, which both methods alleviated similarly. Future work could investigate cluttered scenes, in which mere proximity of the center of gaze to different objects might be insufficient.
@@
The nature of the BBT allowed P1 in Study 1 to achieve very fast transfers through skilled movement of the upper body as a whole. While all participants were informed that they were free to move thus, and in fact, would need to do so in order to reach all blocks, only one was able to use body movement to such an extent and with such skill. Had all adopted this strategy, no results of interest might have been obtained. This motivates future studies based on a more spatially complex task (such as the PHAM [69, 28] or the pasta box task [70]).
All our participants had had at least occasional prior exposure to video games. It was found that users naïve to EMG who also have no experience of video games struggle to use hand gestures as independent controls. This is a limitation of the studies in their current form; it is not clear whether a slow-paced “zeroth” session might address it. For participants like P3-P5 in Study 2, who were able to pass the set criteria for classification but still struggled with control during their first session, there is no doubt that a preliminary session before the first day (as well as a more stringent classification test, such as $75\%$ accuracy in a longer trial [71]) would allow for a fairer comparison between methods.
@@
As participants were aware of the nature of each method, it is difficult to judge whether the minor improvement in Study 1 with D relative to C in success rates may be attributed to participant bias. However, as D occurs before and after C the same number of times in Study 1, learning and fatigue can both be ruled out as explanations. The effect of modeling sequential tasks merits further investigation with a larger dataset.
@@
Six of eight participants selected D as the method they would most prefer if attempting an additional trial. P2 selected C, which is explained by the deterioration of EMG acquisition in the first iteration of D (Table 3). P7 selected B, explaining that, even though it was more difficult than C/D, it made him feel the most satisfied on success, because he felt in control. This is especially interesting in light of the fact that during trials with B, P7 had declared a general weakness in spatial reasoning, which was most challenged by end-effector control. Understandably, P7 alone ranked B above D on the question of agency. P1 ranked A as second in agency and perceived speed, which may be explained by the technical issue in the first iteration of C. Aside from these, all participants in Study 1 ranked C and D as the top two in all three metrics: ease, agency, and perceived speed. The design choice to move the arm along the planned trajectory only as long as EMG input was actively provided seems to have been sufficient for most participants’ perceived sense of control. While these studies do not constitute a thorough analysis of agency or embodiment, which are crucial for prosthesis users, ProACT might serve as a tool for such studies in the future.
@@
In the comparison between A and B, exactly half of the participants preferred A over B and vice-versa. At first glance, this may appear to be a natural variation in preferences. However, examining Tables (35) in conjunction reveals that each participant preferred whichever of these two methods they had performed first; very likely a fatigue effect. It is encouraging that in the light of such an effect, the improvements offered by C and D (which were encountered in alternation, as shown in Table 3) overwhelmed fatigue to keep them at the top of nearly all rankings.
Study 2 indicates the need for further improvements, e.g., feedback about the status of the planner, and a double confirmation for opening the hand to reduce the risk of dropping. Those who faced uncertainty in EMG control also remarked that biofeedback would be helpful in distinguishing it from the uncertainty of planning success.
It is of interest to study the effect of methods which use gaze, whether actively or passively, upon gaze behavior during the task. For example, P1 in Study 2, looked at the upper arm more often in A than in D, the additional visual attention, by his own description, aiding in control. All participants looking at the hand often during D also suggests that they were more unsure of being able to predict its behavior, due to the autonomous motion planning. Following [72], gaze can be developed into a useful tool for evaluating the level of trust/certainty when using methods in the platform.
ProACT can also be used to further improve conventional methods like joint-based control, e.g., by studying which joints are used more heavily in tasks, and informing decisions about which subset of modes to cycle through, as is already done in commercial prostheses, or intelligently select and automate certain joints at certain times. Indeed, some users might prefer such a form of partial autonomy, as compared to task-space motion planning. It is through performing rigorous studies with large and generalizable samples that these questions might be answered; it is our hope that ProACT will enable such investigations.
@@ 5 Conclusion
The present work demonstrates the following:
@@
    Gaze-based intent estimation and low-level autonomy have the potential to improve the performance and comfort of users of high-DoF prosthetic arms in complex manipulation tasks, based on preliminary data from non-amputees.
@@
Due to the open-source nature of much of the latest robotics infrastructure, as well as ProACT itself, it is easy to incorporate more sophisticated methods of intent prediction, and test them in the context of more challenging tasks, a different level of limb loss, or other wearable robotic manipulators. Future work with greater powered samples, as well as with amputee populations, will provide some of the insights necessary for the realization of intelligent whole-arm prostheses.